To prevent attacks of angina, isosorbide mononitrate is given twice daily typically at doses of 20 mg. It is administered at specific times, such as 8 AM and 3 PM, to prevent nitrate tolerance, which causes the drug to be less effective. How many 10-mg tablets are required for 30 days of treatment?
Answer #1

Ordered dose= 20 mg twice daily

Total dose per day= 40 mg(20*2=40)

Available dose= 10mg

Total tablets for one day= 40/10= 4

so, total tablets for 30 days= 30* 4 = 120 tablets.